OpenStudy (jozelynw):

A local snail population grows according to the function g(x) = 200(1.03)2x. Demonstrate the steps to convert g(x) into an equivalent function with only x as the exponent. Then, explain to Iris how the key features of this local snail population compares to the key features of the invasive population. I did the first part, but I need help doing the second part.

OpenStudy (thomas5267):

Note: \[ a^{bc}=\left(a^b\right)^c\\ \text{For example:}\\ 2^{3\times2}=64\\ \left(2^3\right)^2=8^2=64 \]

OpenStudy (jozelynw):

@thomas5267 what exponential property is that?

OpenStudy (jozelynw):

so like this g=200(1.03^2)^x

OpenStudy (jozelynw):

@thomas5267

OpenStudy (thomas5267):

I am unaware of a name for that property, even after a Google search. It just works. To prove this property, you first have to define exponentiation, which is a non-trivial task. I would recommend you to just use it and not worry about it.

OpenStudy (jozelynw):

@nickmeneses1997 and @MathMan77 @mathmate

OpenStudy (anonymous):

I am also not sure of the name but I know that is correct, sorry

OpenStudy (jozelynw):

@MathMan77 what about the second part.

OpenStudy (jozelynw):

@emma.monsterr